when 63 is added to 1 fifth off my number the result is double my number
Answer by anantha(86)

sol:
let the number be x
1 fifth of the number=x/5
63 is added to 1 fifth of the number=63+x/5
double my number=2x
according to the problem
63 is added to 1 fifth of the number=double the number
63+x/5=2x
multiplying by 5 on both sides
63*5+x/5*5=2x*5
315+x=10x
subtract x on both sides
315+x-x=10x-x
315=9x
dividing by 9 on both sides
315/9=9x/9
35=x
x=35
the number=35
verification:
1 fifth of the number=35/5=7
63 is added to 1 fifth of the number=63+7=70
double the number=2*35=70
according to the problem both are equal
x=35 is correct answer